OpenStudy (anonymous):

will fan+medal!!! !!!! PLEASE HELP ME!!!!!!!!!!!!!! Game wardens use experiments to help determine the number of fish in a lake. Suppose 80 fish are caught, tagged and releases back into the lake. Two weeks later 120 fish are caught, 3 of which are found to have tags. Using this information, estimate the number of fish in the lake. a. 3,200 b. 600 c. 120 d. 280

OpenStudy (campbell_st):

this is called capture, re-capture and used ratios to estimate populations let p = the population then \[\frac{p}{80} = \frac{120}{3}\]

OpenStudy (anonymous):

so, it would be A?

OpenStudy (campbell_st):

you need to solve for p... the method is \[\frac{populatio}{tagged} = \frac{re-capture}{re-capture~~tagged}\] it assumes the ratio in the recapture is the same across the entire population

OpenStudy (campbell_st):

that's correct